What is the Walmart division break-even in sales dollars?

What is the Target division break-even in sales dollars?

What is the company's overall break-even in sales dollars?

What would be the company's overall net operating income if the company's operated at its two division break-even points?

Total company Walmart division Target division
Sales $572,000 $222,000 $350,000
Variable expenses 178,940 59,940 119,000
Contribution margin 393,060 162,060 231,000
Traceable fixed expenses 252,000 98,000 154,000
Segment margin 141,060 64,060 77,000
Common fixed expense 91,520 35,520 56,000
Net operating income $49,540 $28,540 $21,000

The common fixed expenses have been allocated to the divisions on the basis of sales
Answer #1

Solution a, b and c:

Computation of Contribution margin ratio
Total Company Walmart Division Target Division
Sales $572,000 $222,000 $350,000
Variable Expenses $178,940 $59,940 $119,000
Contribution margin $393,060 $162,060 $231,000
Contribution margin Ratio (Contribution margin / Sales) 68.72% 73.00% 66.00%
Break-even in sales Dollars (Fixed expenses / Contribution margin Ratio) ($252000+$91520)/68.72% = $98000/73% = $154000/66% =
$499,906.99 $134,246.58 $233,333.33

Solution d:

If the company's operated at its two division break-even points, then the company will not be able to recover Common fixed expenses.

Therefore overall Net operating income will be ($91,520). i.e. loss of $91,520.